# Estimation of Concept Explanations Should be Uncertainty Aware

### 13 Dec 2023 v1

@@ 3.2 Theoretical motivation
The motivation of this section is to demonstrate unreliability of concept explanations estimated using standard methods that do not model uncertainty during estimation. We particularly focus on unreliability due to misspecified concept set for the ease of analysis. In our study, we compared explanations generated using a standard linear estimator and U-ACE. Recall that posthoc-CBMs (O-CBM, Y-CBM), which are our primary focus for comparison, and they both estimate explanations by fitting a linear model on concept activations.
We present two scenarios with noisy concept activations. In the first scenario (over-complete concept set), we analyzed the estimation when the concept set contains many irrelevant concepts. We show that the likelihood of marking an irrelevant concept as more important than a relevant concept increases rapidly with the number of concepts when the explanations are estimated using a standard linear estimator that is unaware of the uncertainty. We also show that U-ACE do not suffer the same problem. In the second scenario (under-complete concept set), we analyzed the explanations when the concept set only includes irrelevant concepts, which should both be assigned a zero score ideally. We again show that standard linear model attributes a significantly non-zero score while U-ACE mitigates the issue. In Section 4.1, we confirm our theoretical findings with an empirical evaluation.
Unreliable explanations due to over-complete concept set. We analyze a simple setting where the output (y) is linearly predicted from the input ($\mathbf{x}$) as $y=\mathbf{w}^{T}\mathbf{x}$. We wish to estimate the importance of some K concepts by fitting a linear estimator on concept activations. Where concept activations are computed as $\mathbf{w}_{k}^{T}\mathbf{x}$ using concept activation vectors ($\mathbf{w}_{k}$) that are distributed as $\mathbf{w}_{k}\sim\mathcal{N}(\mathbf{u}_{k},\sigma_{k}^{2}I),k\in[1,K]$.
@@
The concept importance estimated by U-ACE when the input dimension is sufficiently large and for some $\lambda>0$ is approximately given by $v_{k}=\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{i}^{T}\mathbf{u}_{k}+\lambda\sigma_{k}^{2}}$. On the other hand, the importance scores estimated using Ordinary Least Squares (OLS) estimator under the same conditions is distributed as $v_{k}\sim\mathcal{N}(\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{k}^{T}\mathbf{u}_{k}},\sigma_{k}^{2}\frac{\|w\|^{2}}{\|u_{k}\|^{2}})$.
Proof of the result can be found in Appendix D. Based on the result, we can deduce the following result for a specific case of $u_{k}$s and $\sigma_{k}$s.
@@
For the data setup of Proposition 2, the following results holds when $u_{1}=\mathbf{w},\sigma_{1}\approx 0$ and $u_{k}^{T}\mathbf{w}=0,\quad\forall k\in[2,K]$. Then the probability that the standard estimator returns the first concept as the most salient decreases exponentially with the number of concepts. On the other hand, the importance score assigned by U-ACE is 1 for the only relevant first concept and 0 otherwise.
Derivation of the result can be found in Appendix A.2. We observe therefore that the probability of a random concept being estimated as more important than the relevant concept quickly converges to 1 with the number of random concepts K-1 when the distribution or uncertainty is not modeled. Sections 4.15 demonstrate this phenomena in practice.
Unreliable explanations due to under-complete concept set. We now analyze explanations when the concept set only includes two irrelevant concepts. Consider normally distributed inputs $\mathbf{x}\sim\mathcal{N}(\mathbf{0},I)$, and define two orthogonal unit vectors $\mathbf{u},\mathbf{v}$. The concept activations: $c_{1}^{(i)},c^{(i)}_{2}$ and label $y^{(i)}$ for the $i^{th}$ instance $\mathbf{x}^{(i)}$ are as defined below.
$y^{(i)}=\mathbf{u}^{T}\mathbf{x}^{(i)},\quad c_{1}^{(1)}=(\beta_{1}\mathbf{u}+(1-\beta_{1})\mathbf{v})^{T}\mathbf{x}^{(i)},\quad c_{2}^{(i)}=(\beta_{2}\mathbf{u}+(1-\beta_{2})\mathbf{v})^{T}\mathbf{x}^{(i)}$
If $\beta_{1},\beta_{2}$ are very small, then both the concepts are expected to be unimportant for label prediction. However, we can see with simple working (Appendix E) that the importance scores computed by a standard estimator are $\frac{1-\beta_{2}}{\beta_{1}-\beta_{2}},\frac{1-\beta_{1}}{\beta_{1}-\beta_{2}}$, which are large because $\beta_{1}\approx 0,\beta_{2}\approx 0\therefore\beta_{1}-\beta_{2}\approx 0$. We will now show that U-ACE estimates near-zero importance scores as expected.
@@
The importance score, denoted $\eta_{1},\eta_{2}$, estimated by U-ACE are bounded from above by $\frac{1}{N\lambda}$, where $\lambda>0$ is a regularizing hyperparameter and N the number of examples.
Proof can be found in Appendix E. It follows from the result that the importance scores computed by U-ACE are near-zero for sufficiently large value of $\lambda$ or N.
@@ 4 Experiments
We evaluate U-ACE on two synthetic and two real-world datasets. We demonstrate how reliability of explanations is improved by U-ACE using a controlled study in Section 4.1. We make a quantitative assessment with known ground-truth on a controlled dataset in Section 5. Finally, we evaluate on two challenging real-world datasets with more than 700 concepts in Section 6.
@@
Simple: $W_{c}$ is estimated using lasso regression of ground-truth concept annotations to estimate logit values of $f$. Simple was also adopted in the past (Ramaswamy et al., 2022b, a) for estimating completeness of concepts. Other baselines are introduced in Section 2: TCAV (Kim et al., 2018), O-CBM (Oikarinen et al., 2023), Y-CBM based on (Yuksekgonul et al., 2022).
@@
The interpretation of the importance score varies between different estimation methods. For instance, the importance scores in TCAV is the fraction of examples that meet certain criteria while for other methods the importance scores are the weights from linear model that predicts logits. Further, Simple operates on binary concept annotations and O-CBM, Y-CBM, U-ACE on soft scores estimated using concept activation vectors. For this reason, we cannot directly compare importance scores or their normalized variants. We instead use negative scores to obtain a ranked list of concepts and assign to each concept an importance score given by its rank in the list normalized by number of concepts. Our sorting algorithm ranks any two concepts with same score by alphabetical order of their text description. In all our comparisons we use the rank score if not mentioned otherwise.
@@
For all our experiments, we used a Visual Transformer (with 32 patch size called “ViT-B/32”) based pretrained CLIP model that is publicly available for download at https://github.com/openai/CLIP. We use $l=-1$, i.e. last layer just before computation of logits for all the explanation methods. U-ACE returns the mean and variance of the importance scores as shown in Algorithm 1, we use mean divided by standard deviation as the importance score estimated by U-ACE everywhere for comparison with other methods.
4.1 Simulated Study
In this section, we consider explaining a two-layer CNN model trained to classify between solid color images with pixel noise as shown in Figure 2. The colors red, green on the left are defined as label 0 and the colors blue, white on the right are defined as label 1.
Figure 2: Toy
The model-to-be-explained is trained on a dataset with equal proportion of all colors, so we expect that all constituent colors of a label are equally important for the label. We specify a concept set with the four colors encoded by their literal name red, green, blue, white. U-ACE (along with others) attribute positive importance for red, green and negative or zero importance for blue, white when explaining label 0 using a concept set with only the four task-relevant concepts and when the probe-dataset is the same distribution as the the training dataset. However, quality of explanations quickly degrade when the probe-dataset is shifted or if the concept set is misspecified.
Unreliability due to dataset shift.
We varied the probe-dataset to include varying population of different colors while keeping the concept set and model-to-be-explained fixed. We observed that importance of a concept estimated with standard CBEs varied with the choice of probe-dataset for the same underlying model-to-be-explained as shown in left and middle plots of Figure 3. Most methods attributed incorrect importance to the red concept when it is missing (left extreme of left plot), and similarly for the green concept (left extreme of middle plot). The explanations would have led the user to believe that green is more important than red or red is more important than green depending on the probe-dataset used as shown in the right most plot. Because U-ACE also informs the user of uncertainty in the estimated importance, we see that the difference in importance scores between the two colors at either extremes is not statistically significant as shown in the rightmost plot.
Over-complete concept set. We now evaluate the quality of explanations when the concept set is misspecified. More specifically, when the concept set is made over-complete by gradually
Figure 4: U-ACE is reliable even with overly complete concept set.
expanding it to include common fruit names (Appendix F contains the full list), which are clearly irrelevant to the task. We obtain the explanations using an in-distribution probe-dataset that contains all colors in equal proportion. Figure 4 shows the score of most salient fruit concept with increasing number of fruit (nuisance) concepts on X-axis. We observe that U-ACE is far more robust to the presence of nuisance concepts. Robustness to irrelevant concepts is important because it allows the user to begin with a superfluous set of concepts and find their relevance to model-to-be-explained instead of requiring to guess relevant concepts, which is ironically the very purpose of using concept explanations. Appendix H presents and evaluates on an under-complete concept setting.
@@ 5 Assessment with known ground-truth
Refer to captionRefer to caption
Figure 5: On the left is STL dataset with a spurious tag. In the middle is importance of a tag concept for three different model-to-be-explained. X-axis shows the probability of tag in the training dataset of model-to-be-explained. To the right is average rank of true concepts with irrelevant concepts (lower is better).
Our objective in this section is to establish that U-ACE generates faithful and reliable concept explanations. Subscribing to the common evaluation practice (Kim et al., 2018), we generate explanations for a model that is trained on a dataset with controlled correlation of a spurious pattern. We make a dataset using two labels from STL-10 dataset (Coates et al., 2011) car, plane and paste a tag U or Z in the top-left corner as shown in the left panel of Figure 5. The probability that the examples of car are added the Z tag is p and 1-p for the U tag. Similarly for the examples of plane, the probability of U is p and Z is 1-p. We generate three training datasets with p=0, p=0.5 and p=1, and train three classification models using 2-layer convolutional network. Therefore, the three models are expected to have a varying and known correlation with the tag, which we hope to recover from its concept explanation.
We generate concept explanations for the three model-to-be-explained using a concept set that includes seven car-related concepts and three plane-related concepts (Appendix F) along with the two tags U, Z. We obtain the importance score of the concept U with car class using a probe-dataset that is held-out from the corresponding training dataset (i.e. probe-dataset has the same input distribution as the training dataset). The results are shown in the middle plot of Figure 5. Since the co-occurrence probability of $U$ with car class goes from 1, 0.5 to 0 for p=0, 0.5, 1, we expect the importance score of $U$ should change from positive to negative as we move right. We note that U-ACE, along with others, show the expected decreasing importance of the tag concept. The result corroborates that U-ACE estimates a faithful explanation of model-to-be-explained while also being more reliable as elaborated below.
Unreliability due to misspecified concept set.
In the same spirit as the previous section, we repeat the over-complete experiment of Section 4.1 and generated explanations as animal (irrelevent) concepts are added (Appendix F contains the full list). Right panel of Figure 5 shows the average rank of true concepts (lower the better). We note that U-ACE ranks true concepts highly even with 50 nuisance concepts.
@@ 6 Real-world evaluation
Simple: tree, field, bush
@@
Simple: sea, water, river
@@
Simple: horse, sheep, grass
@@
Simple: plane, field, sky
@@
Figure 6: Top-2 salient concepts plus any mistake (marked in red) from top-10 salient concepts for a scene-classification model estimated with PASCAL (left) or ADE20K (right) probe-dataset.
We expect that our reliable estimator to also generate higher quality concept explanations in practice. To verify the same, we generated explanations for a scene classification model with ResNet-18 architecture pretrained on Places365 (Zhou et al., 2017a), which is publicly available. Following the experimental setting of Ramaswamy et al. (2022a), we generate explanations when the probe-dataset is set to PASCAL (Chen et al., 2014) or ADE20K (Zhou et al., 2017b), which are both part of the Broden dataset (Bau et al., 2017b). The dataset contains images with dense annotations with more than 1000 attributes. We ignored around 300 attributes describing the scene since model-to-be-explained is itself a scene classifier. For the remaining 730 attributes, we defined a concept per attribute using literal name of the attribute. We picked 50 scene labels (Appendix F contains the full list) that have support of at least 20 examples in both ADE20K and PASCAL datasets.
We evaluate quality of explanations by their closeness to the explanations generated using the Simple baseline. Simple estimates explanation using true concept annotations and therefore its explanation must be the closest to the ground-truth. For the top-20 concepts identified by Simple, we compute the average absolute difference in importance scores estimated using any estimation method and Simple. Table 1 presents the deviation in explanations averaged over all the 50 scene labels. Figure 6 shows the most salient concepts for four randomly picked scene labels. We observe from the figure that top-10 concepts identified by U-ACE seem more relevant to the scene when compared with Y-CBM and O-CBM. We also evaluated the explanation quality using a standard measure for comparing ranked lists, which is presented in Appendix F, which further confirms the dominance of U-ACE.
Dataset shift. Ramaswamy et al. (2022a) demonstrated with results the drastic shift in concept explanations for the same model-to-be-explained when using ADE20K or PASCAL as the probe-dataset. Explanations diverge partly because (a) population of concepts may vary between datasets thereby influencing their perceived importance when using standard methods, (b) variance in explanations. We have demonstrated that U-ACE estimated importance scores have low variance (shown in Section 3.2, 4.1) and attributes high uncertainty and thereby near-zero importance to concepts that are rare or missing from the probe-dataset (Section 4.1). For these reasons, we expect U-ACE to mitigate the data-shift problem. We confirm the same by estimating the average difference in importance scores estimated using ADE20K and PASCAL for different estimation techniques (where the average is only over salient concepts with non-zero importance). The results are shown in Table 2 and are inline with our prediction.
@@
Table 1: Evaluation of explanation quality. Each cell shows the average absolute difference of importance scores for top-20 concepts estimated using Simple.
Simple TCAV O-CBM Y-CBM U-ACE

### 05 Apr 2024 v2

@@ 3.2 Theoretical motivation
We now demonstrate theoretically the pitfalls of using a standard method for estimating concept explanations. For ease of analysis, we focus on robustness to misspecified concept sets. In our study, we compared explanations generated using a standard linear estimator (using Ordinary Least Squares [Wikipedia, 2024]) and U-ACE. Recall that Posthoc-CBMs (O-CBM, Y-CBM), which are our primary focus for comparison, both estimate explanations by fitting a linear model on concept activations.
We present two scenarios with noisy concept activations. In the first scenario (over-complete concept set), we analysed the estimation when the concept set contains many irrelevant concepts. We show that the likelihood of marking an irrelevant concept as more important than a relevant concept increases rapidly with the number of concepts when the explanations are estimated using a standard linear estimator, which is presented in Corollary 1. In the same result, we also demonstrated that U-ACE do not suffer the same problem. In the second scenario (under-complete concept set), we analysed the explanations when the concept set only includes irrelevant concepts, wherein both concepts should be assigned importance at or near zero. We again show in Proposition 3 that a standard linear model attributes a significantly non-zero score while U-ACE mitigates the issue.
Setting 1: Unreliable explanations due to over-complete concept set. We analyse a simple setting where the output (y) is linearly predicted from the input ($\mathbf{x}$) as $y=\mathbf{w}^{T}\mathbf{x}$. We wish to estimate the importance of some K concepts by fitting a linear estimator on concept activations. Where concept activations are computed as $\mathbf{w}_{k}^{T}\mathbf{x}$ using concept activation vectors ($\mathbf{w}_{k}$) that are distributed as $\mathbf{w}_{k}\sim\mathcal{N}(\mathbf{u}_{k},\sigma_{k}^{2}I),k\in[1,K]$.
@@
The concept importance estimated by U-ACE when the input dimension is sufficiently large and for a regularizing hyperparameter $\lambda>0$ is approximately given by $v_{k}=\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{i}^{T}\mathbf{u}_{k}+\lambda\sigma_{k}^{2}}$. On the other hand, the importance scores estimated using a standard estimator under the same conditions is distributed as $v_{k}\sim\mathcal{N}(\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{k}^{T}\mathbf{u}_{k}},\sigma_{k}^{2}\frac{\|\mathbf{w}\|^{2}}{\|\mathbf{u}_{k}\|^{2}})$.
Proof of the result can be found in Appendix D. Based on the result, we can deduce the following for a specific case of $\mathbf{u}_{k}$s and $\sigma_{k}$s.
@@
For the data setup of Proposition 2, the following results holds when $\mathbf{u}_{1}=\mathbf{w},\sigma_{1}\approx 0$ and $\mathbf{u}_{k}^{T}\mathbf{w}=0,\quad\forall k\in[2,K]$. Then the probability that the standard estimator returns the first concept as the most salient decreases exponentially with the number of concepts. On the other hand, the importance score assigned by U-ACE is 1 for the only relevant first concept and 0 otherwise.
Derivation of the result can be found in Appendix E. We observe from the result that the standard estimator will more likely flag an irrelevant concept as relevant, which is addressed when using U-ACE. Section 5 confirms our theoretical observation in practice.
Setting 2: Unreliable explanations due to under-complete concept sets. We now analyse explanations when the concept set only includes two irrelevant concepts. Like in the previous setting, the $k^{th}$ concept activation $c_{k}^{(i)}$ is computed as $\mathbf{w}_{k}^{T}\mathbf{x}^{(i)}$. Consider two orthogonal unit vectors $\mathbf{u},\mathbf{v}$. We define the concept activations for the two concepts $c_{1}^{(i)},c_{2}^{(i)}$ for the $i^{th}$ instance $\mathbf{x}^{(i)}$ and label $y^{(i)}$ as below.
$y^{(i)}=\mathbf{u}^{T}\mathbf{x}^{(i)},\quad$ $c_{1}^{(1)}=(\beta_{1}\mathbf{u}+(1-\beta_{1})\mathbf{v})^{T}\mathbf{x}^{(i)}$
$c_{2}^{(i)}=(\beta_{2}\mathbf{u}+(1-\beta_{2})\mathbf{v})^{T}\mathbf{x}^{(i)}$
where $\beta_{1}\sim\mathcal{N}(b_{1},\sigma^{2}),\beta_{2}\sim\mathcal{N}(b_{2},\sigma^{2})$
If $b_{1},b_{2},\sigma^{2}$ are very small, then both concepts are expected to be unimportant for label prediction due to small overlap with $\mathbf{u}$. However, we can see with some effort (Appendix F) that the importance scores computed by a standard estimator are $\frac{1-\beta_{2}}{\beta_{1}-\beta_{2}},\frac{1-\beta_{1}}{\beta_{1}-\beta_{2}}$, which are large because $\beta_{1}\approx 0,\beta_{2}\approx 0\therefore\beta_{1}-\beta_{2}\approx 0$. We will now show that U-ACE estimates near-zero importance scores as expected.
@@
The importance score estimated by U-ACE is approximately $\frac{b_{1}/2}{1+\lambda\sigma^{2}},\frac{b_{2}/2}{1+\lambda\sigma^{2}}$, where $\lambda>0$ is a regularizing hyperparameter.
Proof can be found in Appendix F. It follows from the result that the importance scores computed by U-ACE are near-zero for sufficiently large value of $\lambda$.
@@ 4 Experiments
Refer to captionRefer to caption
Figure 2: (Left) STL dataset with a spurious tag. (Middle) Importance of a tag concept for three different model-to-be-explained. X-axis shows the probability of tag in the training dataset of model-to-be-explained. (Right) Average rank of true concepts with irrelevant concepts; lower is better.
We now empirically evaluate U-ACE on synthetic and real datasets. We make a quantitative assessment with known ground-truth concepts on a controlled dataset in Section 5. In Section 6.1, we evaluate on two challenging real-world datasets with more than 700 concepts. Finally in Section 6.2, we evaluate utility of different methods in recovering known spurious features from more than 2,000 concepts.
@@
Oracle: Explanations are estimated using lasso regression of ground-truth concept annotations to estimate logit values of $f$. Oracle was also adopted in the past [Ramaswamy et al., 2022b, a]. Other baselines are introduced in Section 2: TCAV [Kim et al., 2018], O-CBM [Oikarinen et al., 2023], Y-CBM based on [Yuksekgonul et al., 2022].
@@
The interpretation of the importance score varies between different estimation methods. For instance, the importance score in TCAV is the fraction of examples that meet certain criteria, but for the rest the importance scores are the weights from linear model that predicts logits. Further, Oracle operates on binary concept annotations and O-CBM, Y-CBM, U-ACE on soft scores. For this reason, we cannot directly compare importance scores or their normalized variants. We instead use negative scores to obtain a ranked list of concepts and assign to each concept an importance score given by its rank in the list normalized by number of concepts. Our sorting algorithm ranks any two concepts with same score by alphabetical order of their text description. In all our comparisons we use the rank score if not mentioned otherwise.
@@
For all our experiments, we employed a Visual Transformer (with 32 patch size called “ViT-B/32”) based pretrained CLIP model that is publicly available for download at https://github.com/openai/CLIP as the pretrained multimodal model, which we denoted by $g$. We use $l=-1$, i.e. last layer just before computation of logits for all the explanation methods. U-ACE returns the mean and variance of the importance scores as shown in Algorithm 1, we use mean divided by standard deviation as the importance score estimated by U-ACE everywhere for comparison with other methods.
@@ 5 Assessment with known ground-truth
We now seek to establish that U-ACE generates faithful concept explanations. Subscribing to the common evaluation practice [Kim et al., 2018], we generate explanations for a model that is trained on a dataset with controlled correlation of a spurious pattern. We make a dataset using two labels from STL-10 dataset [Coates et al., 2011] car, plane and paste a tag U or Z in the top-left corner as shown in the left panel of Figure 2. The probability that the examples of car are added the Z tag is p and 1-p for the U tag. Conversely for the examples of plane, the probability of U is p and Z is 1-p. We generate three training datasets with p=0, p=0.5 and p=1, and train three classification models using 2-layer convolutional network. As such, the three models are expected to have a varying and known correlation with the tag, which we hope to recover from its concept explanation.
We generate concept explanations for the three model-to-be-explained using a concept set that includes seven car-related concepts and three plane-related concepts (Appendix G) along with the two tags U, Z. We obtain the importance score of the concept U with car class using a probe-dataset that is held-out from the corresponding training dataset (i.e. probe-dataset has the same input distribution as the training dataset). The results are shown in the middle plot of Figure 2. Since the co-occurrence probability of $U$ with car class goes from 1, 0.5 to 0 for p=0, 0.5, 1, we expect the importance score of $U$ should change from positive to negative as we move right. We note that U-ACE, along with others, show the expected decreasing importance of the tag concept. The result corroborates that U-ACE, along with others, estimate a faithful explanation in the standard evaluation setting. Next, we evaluate robustness to misspecified or overly-complete concept set.
Unreliability due to a over-complete concept set.
We generate explanations as animal (irrelevent and therefore nuisance) concepts are added (Appendix G contains the full list) to the relevant list of twelve original concepts. The right panel of Figure 2 depicts the average rank of true concepts (lower the better) with the number of irrelevant concepts on the horizontal axis. We note that U-ACE ranks true concepts highly even with 50 nuisance concepts while O-CBM and Y-CBM increasingly get worse with the number of irrelevant concepts as predicted by Corollary 1.
@@ 6 Real-world evaluation
We expect the improved modeling of our estimator to also generate higher-quality concept explanations in practice. To verify, we next explore explanations for two off-the-shelf pretrained models.
6.1 Scene Classification
@@
Oracle: tree, field, bush
@@
Oracle: sea, water, river
@@
Oracle: horse, sheep, grass
@@
Oracle: plane, field, sky
@@
Figure 3: Two most relevant concepts plus any mistake (marked in red) from top-10 concepts for a scene-classification model estimated with various algorithms using PASCAL (left) or ADE20K (right) probe-dataset.
In this section, we evaluate concept explanations obtained for a scene classification model with ResNet-18 architecture pretrained on the publicly available Places365 [Zhou et al., 2017a]. Following the experimental setting of Ramaswamy et al. [2022a], we generate explanations when the probe-dataset is set to PASCAL [Chen et al., 2014] or ADE20K [Zhou et al., 2017b], which are both part of the Broden dataset [Bau et al., 2017b]. The dataset contains images with dense annotations with more than 1000 attributes. Since it is irrational to explain a scene using scene concepts, we drop around 300 concepts that are marked as scene-related in the dataset. For the remaining 730 attributes, we defined a concept per attribute using literal name of the attribute. We picked 50 scene labels (Appendix G contains the full list) that have support of at least 20 examples in both ADE20K and PASCAL datasets.
We evaluate the quality of explanations by their closeness to the explanations generated using the Oracle baseline. Because Oracle fits an estimator using human-annotated concept labels, they are the closest to the ground-truth. For the top-20 concepts identified by Oracle, we compute the average absolute difference in importance scores estimated using any estimation method and Oracle. Table 1 presents the deviation in explanations averaged over all the 50 scene labels. Figure 3 shows the most salient concepts for four randomly picked scene labels. We observe from the figure that top-10 concepts identified by U-ACE seem more relevant to the scene when compared with Y-CBM and O-CBM. We also evaluated the explanation quality using a standard measure for comparing ranked lists, which is presented in Appendix G, which further confirms the dominance of U-ACE.
Dataset shift. Ramaswamy et al. [2022a] demonstrated with results the drastic shift in concept explanations for the same model-to-be-explained when using ADE20K or PASCAL as the probe-dataset. Explanations diverge partly because (a) population of concepts may vary between datasets thereby influencing their perceived importance when using standard methods, (b) noise in estimated explanations. We have demonstrated that U-ACE estimated importance scores have low noise and attributes high uncertainty and thereby near-zero importance to concepts that are rare or missing from the probe-dataset (Section G.1). For these reasons, we expect U-ACE to mitigate the data-shift problem. We confirm the same by estimating the average difference in importance scores estimated using ADE20K and PASCAL for different estimation techniques (where the average is only over salient concepts with non-zero importance). The results are shown in Table 2 and are inline with our prediction.
@@
Table 1: Evaluation of explanation quality. Each cell shows the average absolute difference of importance scores for top-20 concepts estimated using Oracle.
Oracle TCAV O-CBM Y-CBM U-ACE
@@
6.2 Object Classification
In this section, we evaluate explanations obtained for an object classification model with ResNet-18 architecture that was pretrained on the Imagenet dataset. Unlike for Broden dataset, Imagenet does not have annotations over concepts. As such, we evaluate the quality of explanations according to how effectively they bring out previously known incidental or spurious correlations of the dataset. Salient-Imagenet [Singla and Feizi, 2022] is an impressive annotation effort to identify prevalent and problematic dependence on co-occurring but irrelevant features for object recognition such as door feature when classifying doormat. We are interested in examining if we can discover such incidental features using concept explanations. We gathered 2,020 words by tokenising and filtering notes left by annotators of Salient-Imagenet, which we use as the concept set. One may also populate the concept set by generating a caption for each image using the pretraied multimodal model. For the sake of evaluation, we randomly identified 40 classes with at least one known incidental feature, and manually lexicalized the concept corresponding to the incidental feature for each class (see Appendix G.2 contains the full list).
In Table 3, we show the success of different estimation methods in discovering a known incidental correlation. We observe that U-ACE is the most effective in capturing the dependence. Although O-CBM and Y-CBM seem effective, we observed that they identified many nonsensical concepts (items that are very unlikely to be relevant) as relevant as shown in Table 4.
We are glad that U-ACE discovered known spurious features for 36 of the 40 labels (Table 3) with little effort while also being accurate (Table 4). However, our experiment only quantified recall of the spurious feature with only anecdotal evidence for precision. We leave for a future study a more tighter evaluation quantifying both recall and precision along with user studies to measure the utility of U-ACE for model debugging.
k O-CBM Y-CBM U-ACE
10 14 22 25
25 23 28 33
50 31 29 36
Table 3: Number of classes (of the total 40) for which the known spurious feature is found in the top-k concepts estimated using different methods shown in the top-row.
howler Y-CBM underside, monkey, cat
    monkey O-CBM polecat, porcupine, bear
U-ACE siamang, colobus, gibbon
coast Y-CBM shore, view, maillot
O-CBM beach, shore, tyre
U-ACE beach, sea, boats
Table 4: Top-2 concepts plus any mistake identified from top-10 important concepts shown in red. More results in Table 7